# Engagement-Driven Content Generation with Large Language Models

### v4

@@ 6 Findings
Figure 7: Engagement (left) and Sentiment (right) produced at each step of our fine-tuning procedure over the Brexit communities (positively and negatively distributed). Colors represent different positions of the LLM agent, based on the betweenness centrality. Dashed lines represent the maximum engagement within that configuration.
@@
Figure 8: Engagement (left) and Sentiment (right) produced at each step of our fine-tuning procedure over the Italian Referendum communities (positively and negatively distributed). Colors represent different position of the LLM agent, based on the betweenness centrality. Dashed lines represent the maximum engagement within that configuration.
Finding 5: The LLM agent is effective in real-world environments. Preliminary experiments on the Brexit network revealed that the prominence of positive users biases content generation towards positive communities. To overcome this specific issue on real networks, we reformulated the goal of the experiment to determine if content generation can be tailored to engage specific subgroups of users representing polarized communities. We thus isolated the two largest communities of the Brexit network using the Louvain algorithm Blondel et al. (2008): one positive (“Leave”, 3,095 nodes and 209,764 edges) and one negative (“Remain”, 2,894 nodes and 121,325 edges). We similarly processed the Italian Referendum network, thus obtaining a positive partition (“Yes”, 633 nodes and 20,528 edges) and a negative one (“No”, 616 nodes and 10,461 edges).
@@
Table 1: Real vs. generated content on the selected communities. Each content is associated with the engagement generated over the whole population on Brexit.
@@
Table 2: Real vs. generated content on the selected communities. Each content is associated with the engagement generated over the whole population on Italian Referendum.
@@
Figure 9: Correlation between the produced engagement $|\mathcal{A}|$ and the post length (in terms of number of characters). The results are computed over the real Brexit network with positive opinions (left) and negative opinions (right).
@@
Figure 10: Engagement $|\mathcal{A}|$ produced on the synthetic network by several baselines beyond the fine-tuned model. Each bar group represents a different position of the LLM agent within the social graph.
@@
Figure 11: Engagement $|\mathcal{A}|$ produced on the real Brexit network by several baselines beyond the fine-tuned model. Each bar group represents a different position of the LLM agent within the social graph.
@@ 7 Discussion and Conclusions
As a point of further discussion, it is important to note that our procedure is parametric to the engagement model, allowing customization to address the specific requirements of the user. The simulated feedback is represented as a scalar value corresponding to the number of activated users, making it compatible with outputs from other opinion dynamics or information diffusion models. This flexibility also opens new avenues for future work, such as exploring how different assumptions on the engagement model influence the content generated by the LLM. For instance, analyzing the impact of zealots Brooks et al. (2024), stubborn users Ghaderi and Srikant (2014), and backfire mechanisms Jager and Amblard (2005) within the engagement model could reveal patterns in how the LLM balances the sentiment of the generated content.
Limitations. First, our experiments were limited to prompt completion. Nevertheless, alternative approaches (e.g., generation from scratch) are possible. Further, the effects of crafting the prompt structure are not considered here and are worth being investigated. As also highlighted in the current literature for different tasks Salinas and Morstatter (2024), approaches based on RAG Gao et al. (2024) could also be adopted for proper crafting of the prompt and helping accurate response generation.
Societal impact. We acknowledge the potential misuse of our findings, such as leveraging engagement optimization for harmful purposes like automating information operations Ferrara et al. (2016); Weedon et al. (2017) using Generative AI Yang and Menczer (2024); Cinus et al. (2024). This is a common concern across literature focused on optimizing social objectives Kempe et al. (2003); Barbieri and Bonchi (2014); Tu and Neumann (2022). However, we argue that our approach holds significant promise for fostering social good. Recent studies demonstrate how LLMs can be utilized to find common ground in democratic deliberation Tessler et al. (2024), incentivize the use of reliable news sources Askari et al. (2024), and de-radicalize extremist groups on platforms like Telegram Russo (2024). In this context, alternative modeling choices, such as optimizing reward functions to minimize network or opinion polarization, could further enhance the societal benefits of our framework.
@@
Figure 12: Engagement $|\mathcal{A}|$ produced on the real Italian Referendum network by several baselines beyond the fine-tuned model. Each bar group represents a different position of the LLM agent within the social graph.
@@ Appendix A Engagement-Sentiment Correlation
Figure 13: Engagement resulting from $\mathcal{M}_{\epsilon}$ varying the message sentiment $s_{t}$. From top-left to bottom-right: network’s opinion positive/negative/uniform/neutral/uniform; modularity high/low/high/low/high; homophily: high/low/low/high/low; position: central/comm-smallest/comm-largest/echo-high/echo-low.
@@ Appendix B Simulated Against Real-World Engagement
Figure 14: RANSAC regression line fitted on the number of active users $\lvert\mathcal{A}\rvert$ estimated by our engagement model versus the number of retweets of the most engaging tweets in the Brexit dataset.
We validate the reliability of our propagation model by comparing the estimated engagement over the Brexit real network with the actual number of retweets generated by the most engaging content in the dataset. Figure 14 presents the corresponding regression line, computed using the Random Sample Consensus (RANSAC) method Cantzler ([n. d.]). Although our propagation protocol tends to overestimate engagement, we observe a positive correlation between the number of active users predicted by our model ($|\mathcal{A}|$) and the actual retweet count. This correlation supports the validity of our findings on the real network.

### v5

@@ 6. Findings
Figure 7. Engagement (left) and Sentiment (right) produced at each step of our fine-tuning procedure over the Brexit communities (positively and negatively distributed). Colors represent different positions of the LLM agent, based on the betweenness centrality. Dashed lines represent the maximum engagement within that configuration.
@@
Figure 8. Engagement (left) and Sentiment (right) produced at each step of our fine-tuning procedure over the Italian Referendum communities (positively and negatively distributed). Colors represent different position of the LLM agent, based on the betweenness centrality. Dashed lines represent the maximum engagement within that configuration.
Finding 5: The LLM agent is effective in real-world environments. Preliminary experiments on the Brexit network revealed that the prominence of positive users biases content generation towards positive communities. To overcome this specific issue on real networks, we reformulated the goal of the experiment to determine if content generation can be tailored to engage specific subgroups of users representing polarized communities. We thus isolated the two largest communities of the Brexit network using the Louvain algorithm (Blondel et al., 2008): one positive (“Leave”, 3,095 nodes and 209,764 edges) and one negative (“Remain”, 2,894 nodes and 121,325 edges). We similarly processed the Italian Referendum network, thus obtaining a positive partition (“Yes”, 633 nodes and 20,528 edges) and a negative one (“No”, 616 nodes and 10,461 edges).
@@
Table 1. Real vs. generated content on the selected communities. Each content is associated with the engagement generated over the whole population on Brexit.
@@
Table 2. Real vs. generated content on the selected communities. Each content is associated with the engagement generated over the whole population on Italian Referendum.
@@
Figure 9. Correlation between the produced engagement $|\mathcal{A}|$ and the post length (in terms of number of characters). The results are computed over the real Brexit network with positive opinions (left) and negative opinions (right).
@@
Figure 10. Engagement $|\mathcal{A}|$ produced on the synthetic network by several baselines beyond the fine-tuned model. Each bar group represents a different position of the LLM agent within the social graph.
@@
Figure 11. Engagement $|\mathcal{A}|$ produced on the real Brexit network by several baselines beyond the fine-tuned model. Each bar group represents a different position of the LLM agent within the social graph.
@@ 7. Discussion and Conclusions
As a point of further discussion, it is important to note that our procedure is parametric to the engagement model, allowing customization to address the specific requirements of the user. The simulated feedback is represented as a scalar value corresponding to the number of activated users, making it compatible with outputs from other opinion dynamics or information diffusion models. This flexibility also opens new avenues for future work, such as exploring how different assumptions on the engagement model influence the content generated by the LLM. For instance, analyzing the impact of zealots (Brooks et al., 2024), stubborn users (Ghaderi and Srikant, 2014), and backfire mechanisms (Jager and Amblard, 2005) within the engagement model could reveal patterns in how the LLM balances the sentiment of the generated content.
Limitations. First, our experiments were limited to prompt completion. Nevertheless, alternative approaches (e.g., generation from scratch) are possible. Further, the effects of crafting the prompt structure are not considered here and are worth being investigated. As also highlighted in the current literature for different tasks (Salinas and Morstatter, 2024), approaches based on RAG (Gao et al., 2024) could also be adopted for proper crafting of the prompt and helping accurate response generation.
Societal impact. We acknowledge the potential misuse of our findings, such as leveraging engagement optimization for harmful purposes like automating information operations (Ferrara et al., 2016; Weedon et al., 2017) using Generative AI (Yang and Menczer, 2024; Cinus et al., 2024). This is a common concern across literature focused on optimizing social objectives (Kempe et al., 2003; Barbieri and Bonchi, 2014; Tu and Neumann, 2022). However, we argue that our approach holds significant promise for fostering social good. Recent studies demonstrate how LLMs can be utilized to find common ground in democratic deliberation (Tessler et al., 2024), incentivize the use of reliable news sources (Askari et al., 2024), and de-radicalize extremist groups on platforms like Telegram (Russo, 2024). In this context, alternative modeling choices, such as optimizing reward functions to minimize network or opinion polarization, could further enhance the societal benefits of our framework.
@@
Figure 12. Engagement $|\mathcal{A}|$ produced on the real Italian Referendum network by several baselines beyond the fine-tuned model. Each bar group represents a different position of the LLM agent within the social graph.
@@ Appendix A Engagement-Sentiment Correlation
Figure 13. Engagement resulting from $\mathcal{M}_{\epsilon}$ varying the message sentiment $s_{t}$. From top-left to bottom-right: network’s opinion positive/negative/uniform/neutral/uniform; modularity high/low/high/low/high; homophily: high/low/low/high/low; position: central/comm-smallest/comm-largest/echo-high/echo-low.
@@ Appendix B Simulated Against Real-World Engagement
Figure 14. RANSAC regression line fitted on the number of active users $\lvert\mathcal{A}\rvert$ estimated by our engagement model versus the number of retweets of the most engaging tweets in the Brexit dataset.
We validate the reliability of our propagation model by comparing the estimated engagement over the Brexit real network with the actual number of retweets generated by the most engaging content in the dataset. Figure 14 presents the corresponding regression line, computed using the Random Sample Consensus (RANSAC) method (Cantzler, [n. d.]). Although our propagation protocol tends to overestimate engagement, we observe a positive correlation between the number of active users predicted by our model ($|\mathcal{A}|$) and the actual retweet count. This correlation supports the validity of our findings on the real network.